# Online Estimation with Rolling Validation: Adaptive Nonparametric Estimation with Streaming Data

### v3

@@ 4 Consistency of Rolling Validation, Fixed Number of Candidates
4.1 Sketch of the Analysis and some Technical Results
The wRV statistics is an unbiased estimator of the estimators’ accumulative population risk (14). Theorem 4.1 assumed a “gap” between the rolling population risks for $g_{i}$ and $h_{i}$. To ensure statistical consistency of wRV, we need to establish some concentration results of the sample RV statistics on their individual means. Specifically, the fluctuation of the sample quantities should be of a smaller order than the magnitude of the population risk gap. Consequently, the realized trajectories of $RV(\{g_{i}\})$ and $RV(\{h_{i}\})$ will eventually separate. To elucidate the concentration properties, we rewrite the difference between sample wRV and its population mean as follows. Define $u_{i}=i^{\xi}\left\{\hat{f}_{i-1}\left(X_{i}\right)-f_{0}\left(X_{i}\right)\right\}^{2}$:
$\sum_{i=1}^{n}i^{\xi}\{\hat{f}_{i-1}(X_{i})-Y_{i}\}^{2}-i^{\xi}\epsilon_{i}^{2}-\sum_{i=1}^{n}i^{\xi}E[\{\hat{f}_{i-1}(X)-f_{0}(X)\}^{2}]$ (20)
$=$ $\sum_{i=1}^{n}u_{i}-2i^{\xi}\epsilon_{i}\{\hat{f}_{i-1}(X_{i})-f_{0}(X_{i})\}-\sum_{i=1}^{n}E[u_{i}]$
$=$ $\left(\sum_{i=1}^{n}u_{i}-E[u_{i}|F^{i-1}]\right)+\left(\sum_{i=1}^{n}E[u_{i}|F^{i-1}]-E[u_{i}]\right)$
@@ 4.1 Sketch of the Analysis and some Technical Results
Assume IID sampling scheme 4.1 and $\{\hat{f}_{i}\}$ satisfying 4.4 with constant $a$, we have
$\lim_{n\rightarrow\infty}\operatorname{pr}\left(\left|\sum_{i=1}^{n}i^{\xi}\epsilon_{i}\left\{\hat{f}_{i-1}\left(X_{i}\right)-f_{0}\left(X_{i}\right)\right\}\right|\geq c_{n}\right)=0$
@@
Assume IID sampling scheme 4.1 and $\{\hat{f}_{i}\}$ satisfying 4.4 with constants $a$, we have
$\lim_{n\rightarrow\infty}\operatorname{pr}\left(\left|\sum_{i=1}^{n}u_{i}-E\left[u_{i}\mid F^{i-1}\right]\right|\geq c_{n}\right)=0$
@@
$\lim_{n\rightarrow\infty}\operatorname{pr}\left(\left|\sum_{i=1}^{n}E\left[u_{i}\mid F^{i-1}\right]-E\left[u_{i}\right]\right|\geq c_{n}\right)=0$ (21)
@@
The proof of Lemma 4.4 is presented in Appendix D.1. We also provide examples in which ${\rm Var}(\sum_{i=1}^{n}E\left[u_{i}\mid F^{i-1}\right]-E\left[u_{i}\right])$ can be explicitly characterized, which implies the requirement of $c_{n}$ in (22) cannot be further relaxed to achieve the desired concentration property (21) in general. See Section D.2 for the derivation and numerical verification.
@@
A typical $c_{n}$ that we are interested in is $c_{n}=n^{1-a+\xi}$—this is the order of $\sum_{i=1}^{n}E[u_{i}]$ when $a\in[0,1)$. If we want the deviation $\left|\sum_{i=1}^{n}E[u_{i}|F^{i-1}]-E[u_{i}]\right|$ to be of a smaller order than the mean $\sum_{i=1}^{n}E[u_{i}]$, we need the stability index $b$ to be greater than $(a+1)/2$ (stated in Theorem 4.1). As $a$ getting larger, the convergence rate of the considered estimator gets better, but the stability requirement becomes more stringent.
@@
$\lim_{n\rightarrow\infty}\operatorname{pr}\left(\left|\sum_{i=1}^{n}i^{\xi}\left\{\hat{f}_{i-1}\left(X_{i}\right)-Y_{i}\right\}^{2}-i^{\xi}\epsilon_{i}^{2}-\sum_{i=1}^{n}i^{\xi}E\left\{\hat{f}_{i-1}(X)-f_{0}(X)\right\}^{2}\right|\geq c_{n}\right)=0$
@@ 5 Consistency with Diverging Candidate Cardinality
So far, we have focused on comparing the performance of a fixed finite number of candidate estimator sequences specified by their hyperparameters $\{(\lambda_{i}^{(k)}:i\in\mathbb{N}^{+}),~{}k\in[K]\}$. It is also natural to explore a diverging number of model sequences in online learning settings. As more data is collected, researchers may refine or consider a broader range of hyperparameter sequences and explore more modeling possibilities (e.g., including higher-order interaction between interesting covariates).
@@
For each element $\lambda_{i+1}^{(k)}$ in $\Lambda_{i+1}$, we define its representer in $\Lambda_{i}$ as $\lambda_{i-}^{(k)}$. Typically, the representer is the element in $\Lambda_{i}$ that is most “similar” to $\lambda_{i+1}^{(k)}$. If $\Lambda_{i}\subset\Lambda_{i+1}$ then for $k\leq K_{i}$ the representer is simply $\lambda_{i-}^{(k)}=\lambda_{i}^{(k)}$. For new hyperparameters $\lambda_{i+1}^{(k)}\notin\Lambda_{i}$, we find their representers in $\Lambda_{i}$ by minimizing a mathematical distance, or by identifying conceptual relationships between $\Lambda_{i}$ and $\Lambda_{i+1}$. For example, some elements in $\Lambda_{i+1}$ are a refinement of earlier hyperparameters. Once the representer mapping is established, we can map each element in $G_{i+1}$ to one in $G_{i}$, assigning all $\hat{f}_{i+1}^{(k)},k\in[K_{i+1}]$ a previous value $\hat{f}_{i-}^{(k)}$. This process can be traced back to $i=1$, ensuring all sequences of hyperparameters and estimators are well-defined, allowing us to continue our discussion.
In the following two subsections, we will outline the technical conditions necessary for consistency and then formally present our results.
@@ 5.1 Technical Conditions for Consistency
    There exists a constant $C$ such that $\operatorname{pr}\left(|X|\geq Lt\right)\leq Ce^{-t^{\theta}}$ for all $t>0$.
@@
    There exists a constant $c$ such that $(E[X^{q}])^{1/q}\leq cLq^{1/\theta}$ for all $q\geq 1$.
@@ 5.2 Consistency with a Diverging Number of Candidate Sequences
    $\liminf_{n\rightarrow\infty}n^{a_{h,n}}\min_{k\in K_{n}}E[r(h^{(k)}_{n})]=M_{h},$
@@
$\lim_{n\rightarrow\infty}\operatorname{pr}\left[{\rm RV}\left(\left\{g_{i},i\in[n]\right\},\xi\right)<\inf_{k\in[K_{n}]}{\rm RV}\left(\left\{h_{i}^{(k)},i\in[n]\right\},\xi\right)\right]=1,$
@@
(Regarding model number $K_{n}$) In Theorem 5.2, we required the number of candidate models to satisfy $\log K_{n}\lesssim\log n$. In general, one may expect the number of items under comparison (here $K_{n}$) to scale like $\log K_{n}\sim n^{\alpha}$ for some $\alpha>0$ when assuming similar exponential light-tail conditions. This is true if we further require $\liminf_{n}\min_{k\in\left[K_{n}\right]}\left(b_{k}-a_{k}/2-1/2\right)\wedge\left(1-a_{k}\right)>0$. However, the current (U1) condition means to explore the regime where the convergence rates and stability conditions approach the boundary, which can afford at most polynomially many $K_{n}$.
@@ 6 Stability of Estimators
The consistency of wRV model selection relies on three interpretable conditions: IID sampling, MSE convergence, and estimator stability. While the convergence rates of estimators have been extensively studied in the literature, the stability condition is relatively new and deserves more discussion. In this section we evaluate the stability rates for multiple prototypical estimators, providing algorithm-specific conditions for the assumptions in Theorem 4.1 to hold.
@@ 6.1 Batch Sieve Estimators
The concept of stability applies to both online and batch/off-line estimators. We begin our discussion with a batch sieve estimator which is simpler and easier to understand at first reading.
@@
Let $\hat{f}(\cdot)=\hat{f}_{n}(\ \cdot\ ;Z_{n})$, $\hat{f}^{\prime}=\hat{f}_{n}(\ \cdot\ ;Z_{n}^{j})$ be two estimates but are different at the $j$-th training sample. For any $j\in[n]$,
@@
$E\left[\left\{\nabla_{j}\hat{f}_{n}(X)\right\}^{2}\mid F^{j}\right]$ $=n^{-2}E\left[\left(\sum_{k=1}^{J_{n}}\left\{Y_{j}\phi_{k}\left(X_{j}\right)-Y_{j}^{\prime}\phi_{k}\left(X_{j}^{\prime}\right)\right\}\phi_{k}(X)\right)^{2}\mid F^{j}\right]$ (26)
@@
where in step $(I)$ we used $\{\phi_{k}\}$ is orthonormal: $E[\phi_{l}(X)\phi_{k}(X)]=\delta_{lk}$, and in step $(II)$ we assumed the outcome variable and the basis functions are uniformly bounded. We summarize the stability results established above as follows:
@@
Suppose we observe an IID sample $\{(X_{i},Y_{i}),i\in[n]\}$ satisfying Assumption 4.1. We further assume the outcome $Y_{i}$ and basis functions $\{\phi_{k}\}$ are uniformly bounded by $L>0$. Consider orthonormal basis $\{\phi_{k}\}$ w.r.t. $P_{X}$: $E\left[\phi_{l}(X)\phi_{k}(X)\right]=\delta_{lk}$. Then for the estimator defined in (24) and (25) with $J=n^{\alpha}$, we have for any $j\in[n]$:
$n^{2-\alpha}E\left[\left\{\nabla_{j}\hat{f}_{n}(X)\right\}^{2}\mid F^{j}\right]\leq 4L^{4}\text{ almost surely }.$
@@
$E[r(\hat{f})]\lesssim J/n+J^{-2s}Q^{2}.$ (27)
In the case when $\alpha<(2s+1)^{-1}$, the bias term $J^{-2s}Q^{2}$ would dominate the variance term $J/n$ and the convergence rate of $\hat{f}$ will be $n^{-2s\alpha}$. When we use significantly more basis than needed ($\alpha>(2s+1)^{-1}$), the variance term $J/n$ will dominate and the convergence rate behaves like $n^{\alpha-1}$. In order to obtain rate optimal estimators, we need to balance the $J/n$ and the $J^{-2s}Q^{2}$ terms so that they are of the same order. This leads to $J\sim n^{1/(2s+1)}$, with an MSE convergence rate $n^{-2s/(2s+1)}$.
@@ 6.2 Online Parametric SGD
In Section 6.1 we discussed the stability properties of batch sieve estimators. Now we switch back to the online setting. In this section, we consider a linear SGD estimator with update formula
@@
$\hat{f}_{i}(x)=x^{\top}\left(i^{-1}\sum_{k=1}^{i}\beta_{k}\right)\,.$
@@
Assume we observe IID samples $\{X_{i},Y_{i}\}$ satisfying 4.1. We further assume that $\|X_{i}\|\leq R$ and $|Y_{i}-X_{i-1}^{\top}\beta_{i}|\leq M$ almost surely. The minimal eigenvalue $\lambda_{min}(E[XX^{\top}])\geq\underline{\lambda}>0$. When the learning rate $\gamma\leq R^{-1}$, we have, for any $j\in[i]$,
$i^{2}E\left[\left\{\nabla_{j}\hat{f}_{i}\left(X\right)\right\}^{2}\mid F^{j}\right]\leq C(\gamma,\underline{\lambda},M,R)\quad\text{ almost surely },$

### v4

@@ 4 Consistency of Rolling Validation, Fixed Number of Candidates
4.1 Sketch of the Analysis and Key Technical Ingredients
The wRV statistics is an unbiased estimator of the estimators’ cumulative population risk (10). Theorem 4.1 assumed a “gap” between the rolling population risks for $g_{i}$ and $h_{i}$. To ensure statistical consistency of wRV, we need to establish some concentration results of the sample RV statistics on their individual means. Specifically, the fluctuation of the sample quantities should be of a smaller order than the magnitude of the population risk gap. Consequently, the realized trajectories of ${\rm RV}(\{g_{i}\})$ and ${\rm RV}(\{h_{i}\})$ will eventually separate. To elucidate the concentration properties, we rewrite the difference between sample wRV and its population mean as follows. Define $u_{i}=i^{\xi}\left\{\hat{f}_{i-1}\left(X_{i}\right)-f_{0}\left(X_{i}\right)\right\}^{2}$:
$\sum_{i=1}^{n}i^{\xi}\{\hat{f}_{i-1}(X_{i})-Y_{i}\}^{2}-i^{\xi}\epsilon_{i}^{2}-\sum_{i=1}^{n}i^{\xi}\mathbb{E}[\{\hat{f}_{i-1}(X)-f_{0}(X)\}^{2}]$ (15)
$=$ $\sum_{i=1}^{n}u_{i}-2i^{\xi}\epsilon_{i}\{\hat{f}_{i-1}(X_{i})-f_{0}(X_{i})\}-\sum_{i=1}^{n}\mathbb{E}[u_{i}]$
$=$ $\left(\sum_{i=1}^{n}u_{i}-\mathbb{E}[u_{i}|F^{i-1}]\right)+\left(\sum_{i=1}^{n}\mathbb{E}[u_{i}|F^{i-1}]-\mathbb{E}[u_{i}]\right)$
@@ 4.1 Sketch of the Analysis and Key Technical Ingredients
Assume IID sampling scheme 4.1 and $\{\hat{f}_{i}\}$ satisfying 4.4 with a constant $a$, we have
$\lim_{n\rightarrow\infty}\mathbb{P}\left(\left|\sum_{i=1}^{n}i^{\xi}\epsilon_{i}\left\{\hat{f}_{i-1}\left(X_{i}\right)-f_{0}\left(X_{i}\right)\right\}\right|\geq c_{n}\right)=0$
@@
Assume IID sampling scheme 4.1 and $\{\hat{f}_{i}\}$ satisfying 4.4 with a constant $a$, we have
$\lim_{n\rightarrow\infty}\mathbb{P}\left(\left|\sum_{i=1}^{n}u_{i}-\mathbb{E}\left[u_{i}\mid F^{i-1}\right]\right|\geq c_{n}\right)=0$
@@
$\lim_{n\rightarrow\infty}\mathbb{P}\left(\left|\sum_{i=1}^{n}\mathbb{E}\left[u_{i}\mid F^{i-1}\right]-\mathbb{E}\left[u_{i}\right]\right|\geq c_{n}\right)=0$ (16)
@@
The proof of Lemma 4.4 is presented in Appendix D.1. We also provide examples in which ${\rm Var}(\sum_{i=1}^{n}\mathbb{E}\left[u_{i}\mid F^{i-1}\right]-\mathbb{E}\left[u_{i}\right])$ can be explicitly characterized, which implies the requirement of $c_{n}$ in (17) cannot be further relaxed to achieve the desired concentration property (16) in general. See Section D.2 for the derivation and numerical verification.
@@
A typical $c_{n}$ that we are interested in is $c_{n}=n^{1-a+\xi}$—this is the order of $\sum_{i=1}^{n}\mathbb{E}[u_{i}]$ when $a\in[0,1)$. If we want the deviation $\left|\sum_{i=1}^{n}\mathbb{E}[u_{i}|F^{i-1}]-\mathbb{E}[u_{i}]\right|$ to be of a smaller order than the mean $\sum_{i=1}^{n}\mathbb{E}[u_{i}]$, we need the stability index $b$ to be greater than $(a+1)/2$ (stated in Theorem 4.1). As $a$ getting larger, the convergence rate of the considered estimator gets better, but the stability requirement becomes more stringent.
@@
$\lim_{n\rightarrow\infty}\mathbb{P}\left(\left|\sum_{i=1}^{n}i^{\xi}\left\{\hat{f}_{i-1}\left(X_{i}\right)-Y_{i}\right\}^{2}-i^{\xi}\epsilon_{i}^{2}-\sum_{i=1}^{n}i^{\xi}\mathbb{E}[r(\hat{f}_{i-1})]\right|\geq c_{n}\right)=0$
@@ 5 Consistency with Diverging Candidate Cardinality
So far, we have focused on comparing the performance of a fixed number of candidate estimator sequences specified by their hyperparameters $\{(\lambda_{i}^{(k)}:i\in\mathbb{N}^{+}),~{}k\in[K]\}$. It is also natural to explore a diverging number of model sequences in online learning settings. As more data is collected, researchers may refine or consider a broader range of hyperparameter sequences and explore more modeling possibilities (e.g., including higher-order interactions between interesting covariates).
@@
For each element $\lambda_{i+1}^{(k)}$ in $\Lambda_{i+1}$, we define its representer in $\Lambda_{i}$ as $\lambda_{i-}^{(k)}$. Typically, the representer is the element in $\Lambda_{i}$ that is most “similar” to $\lambda_{i+1}^{(k)}$. For $k\leq K_{i}$ the representer is simply $\lambda_{i-}^{(k)}=\lambda_{i}^{(k)}$. For new hyperparameters $\lambda_{i+1}^{(k)}$ with $k>K_{i}$, we find their representers in $\Lambda_{i}$ by minimizing a mathematical distance, or by identifying conceptual relationships between $\Lambda_{i}$ and $\Lambda_{i+1}$. For example, some elements in $\Lambda_{i+1}$ are a refinement of earlier hyperparameters. Once the representer mapping is established, we can map each element in $G_{i+1}$ to one in $G_{i}$, assigning all $\hat{f}_{i+1}^{(k)},k\in[K_{i+1}]$ a previous value $\hat{f}_{i}^{(k)}$. This process can be traced back to $i=1$, ensuring all sequences of hyperparameters and estimators are well-defined, allowing us to continue our discussion.
In the next two subsections, we outline the technical conditions required for consistency and then formally state our results.
@@ 5.1 Technical Conditions for Consistency
    There exists a constant $C$ such that $\mathbb{P}\left(|X|\geq Lt\right)\leq C\exp(-t^{\theta})$ for all $t>0$.
@@
    There exists a constant $c$ such that $(\mathbb{E}[X^{q}])^{1/q}\leq cLq^{1/\theta}$ for all $q\geq 1$.
@@ 5.2 Consistency with a Diverging Number of Candidate Sequences
    $\liminf_{n\rightarrow\infty}n^{a_{h,n}}\min_{k\in K_{n}}\mathbb{E}[r(h^{(k)}_{n})]=M_{h},$
@@
$\lim_{n\rightarrow\infty}\mathbb{P}\left[{\rm RV}\left(\left\{g_{i},i\in[n]\right\},\xi\right)<\inf_{k\in[K_{n}]}{\rm RV}\left(\left\{h_{i}^{(k)},i\in[n]\right\},\xi\right)\right]=1,$
@@
(Regarding model number $K_{n}$) In Theorem 5.2, we required the number of candidate models to satisfy $\log K_{n}\lesssim\log n$. Given the current literature works with light-tail assumptions, one may expect the number of items under comparison (here $K_{n}$) to scale like $\log K_{n}\sim n^{\alpha}$ for some $\alpha>0$ when assuming similar exponential light-tail conditions. This is true if we further require $\liminf_{n}\min_{k\in\left[K_{n}\right]}\left(b_{k}-a_{k}/2-1/2\right)\wedge\left(1-a_{k}\right)>0$. However, the current (U1) condition means to explore the regime where the convergence rates and stability conditions approach the boundary, which can afford at most polynomially many $K_{n}$.
Remark 5.4.
The problem of maintaining a dynamic list of candidate models has been considered in the literature of online functional data analysis. For example, [37] proposed to maintain a list of candidate bandwidths obtained as of weighted average of past candidates and the current candidates. At a high level, this is similar to our proposal of linking a new candidate to the closest one in the previous time step. However, there are two important distinctions between our framework and the setting in [37]. First, the tuning parameter selection in [37] is partially model-based, as it requires pilot estimates of functionals of unknown true and correctly specified model parameters. In contrast, our method is model-free and only requires a loss function to evaluate the predictive quality of the fitted model. Second, the averaging approach taken in [37] requires the candidates to be objects in the same vector space, such as different bandwidths. In contrast, our method is much more flexible as it works for arbitrary collections of model fitting methods. For example, our method can compare local polynomial smoothers with basis expansion estimators, and can even incorporate black-box estimators.
@@ 6 Stability of Estimators
The consistency of wRV model selection relies on three interpretable conditions: IID sampling, MSE convergence, and estimator stability. While the convergence rates of estimators have been extensively studied in the literature, the stability condition is relatively new and deserves more discussion. In this section, we evaluate the stability rates for multiple prototypical estimators, providing algorithm-specific conditions for the assumptions in Theorem 4.1 to hold.
@@ 6.1 Batch Sieve Estimators
The concept of stability applies to both online and batch estimators. We begin our discussion with a batch sieve estimator, which is simpler and easier to understand at first reading.
@@
Let $\hat{f}(\cdot)=\hat{f}_{n}(\ \cdot\ ;\bm{Z}_{n})$, $\hat{f}^{\prime}=\hat{f}_{n}(\ \cdot\ ;\bm{Z}_{n}^{j})$ be two estimates but are different at the $j$-th training sample. For any $j\in[n]$,
@@
$\mathbb{E}\left[\left\{\nabla_{j}\hat{f}_{n}(X)\right\}^{2}\mid F^{j}\right]$ $=n^{-2}\mathbb{E}\left[\left(\sum_{k=1}^{J_{n}}\left\{Y_{j}\phi_{k}\left(X_{j}\right)-Y_{j}^{\prime}\phi_{k}\left(X_{j}^{\prime}\right)\right\}\phi_{k}(X)\right)^{2}\mid F^{j}\right]$ (21)
@@
where in step $(I)$ we used $\{\phi_{k}\}$ is orthonormal: $\mathbb{E}[\phi_{l}(X)\phi_{k}(X)]=\delta_{lk}$, and in step $(II)$ we assumed the outcome variable and the basis functions are uniformly bounded. We summarize the stability results established above as follows:
@@
Suppose we observe an IID sample $\{(X_{i},Y_{i}),i\in[n]\}$ satisfying Assumption 4.1. We further assume the outcome $Y_{i}$ and basis functions $\{\phi_{k}\}$ are uniformly bounded by $L>0$. Consider orthonormal basis $\{\phi_{k}\}$ w.r.t. $P_{X}$: $\mathbb{E}\left[\phi_{l}(X)\phi_{k}(X)\right]=\delta_{lk}$. Then for the estimator defined in (19) and (20) with $J=n^{\alpha}$, we have for any $j\in[n]$:
$n^{2-\alpha}\mathbb{E}\left[\left\{\nabla_{j}\hat{f}_{n}(X)\right\}^{2}\mid F^{j}\right]\leq 4L^{4}\text{ almost surely }.$
@@
$\mathbb{E}[r(\hat{f})]\lesssim J/n+J^{-2s}Q^{2}.$ (22)
In the case when $\alpha<(2s+1)^{-1}$, the bias term $J^{-2s}Q^{2}$ would dominate the variance term $J/n$ and the convergence rate of $\hat{f}$ will be $n^{-2s\alpha}$. When we use significantly more basis than needed ($\alpha>(2s+1)^{-1}$), the variance term $J/n$ will dominate and the convergence rate behaves like $n^{\alpha-1}$. To obtain rate-optimal estimators, we need to balance the $J/n$ and the $J^{-2s}Q^{2}$ terms so that they are of the same order. This leads to $J\sim n^{1/(2s+1)}$, with an MSE convergence rate $n^{-2s/(2s+1)}$.
@@ 6.2 Online Parametric SGD
In Section 6.1 we discussed the stability properties of batch sieve estimators. Now we switch back to the online setting. In this section, we consider a linear SGD estimator with the update formula
@@
$\hat{f}_{i}(x)=x^{\top}\bigg{(}i^{-1}\sum_{k=1}^{i}\beta_{k}\bigg{)}\,.$
@@
Assume we observe IID samples $\{X_{i},Y_{i}\}$ satisfying 4.1. We further assume that $\|X_{i}\|\leq R$ and $|Y_{i}-X_{i-1}^{\top}\beta_{i}|\leq M$ almost surely. The minimal eigenvalue $\lambda_{min}(\mathbb{E}[XX^{\top}])\geq\underline{\lambda}>0$. When the learning rate $\gamma\leq R^{-1}$, we have, for any $j\in[i]$,
$i^{2}\mathbb{E}\left[\left\{\nabla_{j}\hat{f}_{i}\left(X\right)\right\}^{2}\mid F^{j}\right]\leq C(\gamma,\underline{\lambda},M,R)\quad\text{ almost surely },$